### 24 Mar 2026 v4

# Agent Control Protocol ACP v1.17: Admission Control for Agent Actions Technical Specification and Reference Implementation

Abstract
Agent Control Protocol (ACP) is a formal technical specification for governance of autonomous agents in B2B institutional environments. ACP is the admission control layer between agent intent and system state mutation: before any agent action reaches execution, it must pass a cryptographic admission check that validates identity, capability scope, delegation chain, and policy compliance simultaneously.
ACP defines the mechanisms of cryptographic identity, capability-based authorization, deterministic risk evaluation, verifiable chained delegation, transitive revocation, and immutable auditing that a system must implement for autonomous agents to operate under explicit institutional control.
ACP operates as an additional layer on top of RBAC and Zero Trust, without replacing them. It is designed specifically for the problem that neither model solves: governing what an autonomous agent can do, under what conditions, with what limits, and with complete traceability for external auditing—including across organizational boundaries.
The v1.17 specification comprises 38 technical documents organized into five conformance levels (L1–L5). It includes a Go reference implementation of 23 packages covering all L1–L4 capabilities, 73 signed conformance test vectors (Ed25519 + SHA-256) plus 65 unsigned RISK-2.0 vectors, an OpenAPI 3.1.0 specification (18 endpoints), an executable payment-agent demonstration (examples/payment-agent/), and an executable multi-organization interoperability demo. v1.17 adds external verifiability: a TLC-runnable TLA+ formal model with three checked invariants, an ACR-1.0 sequence compliance runner (library mode and HTTP mode), five stateful sequence test vectors covering cooldown, F_anom Rule 3, boundary conditions, privilege jump, and benign flow, and an ACP-SIGN-2.0 HYBRID mode stub for the post-quantum transition. It defines 12 explicitly prohibited behaviors whose presence disqualifies conformance at any level, and the mechanisms for interoperability between institutions.
Specification and implementation: https://github.com/chelof100/acp-framework-en
1  The Problem ACP Solves
Autonomous agents are being deployed in institutional environments without a technical standard to govern their behavior. This is not a tooling problem—it is a protocol problem.
1.1  The Structural Gap
When an autonomous agent makes a decision and executes it, there is a critical moment between both actions. In current models, that moment does not formally exist: the decision and the execution are the same event. The agent decides and acts. There is no intermediate validation. There is no point of intervention. There is no structured record of why the decision was made.
This is acceptable when a human executes that action, because the human bears responsibility and can be questioned. An autonomous agent cannot be questioned. It can only be audited—and only if there is something to audit.
The problem is not whether agents are trustworthy. The problem is that currently no formal technical mechanism exists that allows an institution to demonstrate that its agents operated within authorized limits.
1.2  ACP as Admission Control
The clearest frame for understanding what ACP does is the Kubernetes Admission Controller analogy.
Kubernetes intercepts every API request before it reaches the cluster and runs it through a sequence of admission checks—ValidatingWebhookConfiguration, ResourceQuota enforcement, OPA Gatekeeper policies. If any check fails, the request is rejected before touching cluster state.
ACP applies this pattern to agent actions:
agent intent
[1] Identity check (ACP-AGENT-1.0, ACP-HP-1.0)
| is this agent who they claim to be?
[2] Capability check (ACP-CT-1.0, ACP-DCMA-1.1)
| does the agent hold a valid token for this action?
[3] Policy check (ACP-RISK-2.0, ACP-PSN-1.0)
| is this action within current policy?
[4] ADMIT / DENY / ESCALATE
| (if ADMIT)
[5] Execution token (ACP-EXEC-1.0)
| single-use cryptographic proof of admission
[6] Ledger record (ACP-LEDGER-1.3)
| immutable signed audit entry
system state mutation
The critical difference from Kubernetes: ACP’s admission check operates across institutional boundaries. An agent from Bank A can be admitted by Bank B without Bank B trusting Bank A’s internal infrastructure—the cryptographic proof is self-contained and verifiable with Bank A’s published public key alone.
This admission control framing also clarifies the relationship with related tools:
  • OPA (Open Policy Agent) [10] can serve as the policy evaluation engine inside Step 3. ACP does not replace OPA; it adds the identity and delegation chain layers above it.
  • AWS IAM / Azure RBAC model static role permissions for humans. ACP adds dynamic agent delegation with execution proof.
  • OAuth 2.0 [5] handles API access tokens. ACP extends delegation to multi-agent chains with non-escalation and verifiable provenance.
  • SPIFFE / SPIRE [14] provides cryptographic workload identity. ACP builds on that identity to add capability scoping and governance.
1.3  Why RBAC and Zero Trust Are Insufficient
RBAC and Zero Trust are the predominant control layers in enterprise environments. Both are necessary. Neither solves the problem of governing autonomous agents:
Criterion RBAC Zero Trust ACP
Designed for Human roles Network access Autonomous agents
Cryptographic identity No Partial Yes (Ed25519, mandatory)
Verifiable dynamic delegation No No Yes (chained, auditable)
Decision/execution separation No No Yes (Execution Tokens)
Real-time risk evaluation No Partial Yes (deterministic)
Multi-institutional auditing Non-standard Non-standard Native (signed ledger)
Transitive delegation revocation No No Yes (formal propagation)
B2B interoperability for agents Unstructured Unstructured Central protocol design
Table 1: Comparison of ACP with RBAC and Zero Trust across agent governance criteria.
ACP does not replace RBAC or Zero Trust. It adds a governance layer oriented specifically to autonomous agents that operates above existing controls.
1.4  The Concrete Scenario ACP Prevents
Without ACP: A financial processing agent receives instructions from another agent to execute a transfer. The agent executes the action. If the instruction was legitimate, everything works. If it was compromised, injected, or generated by an unauthorized agent, the transfer occurs anyway. No formal mechanism exists to prevent it, nor is there a technical record that allows reconstructing the authorization chain.
With ACP: The agent requesting the transfer must present a Capability Token cryptographically signed by the issuing institution, demonstrate possession of the associated private key, and the request passes through the risk engine before receiving an Execution Token. The ET is single-use. The entire chain is recorded in the Audit Ledger with an externally verifiable institutional signature.
2  What ACP Is
A formal technical specification—not a framework, not a platform, not a set of best practices. A protocol with precise definitions, formal state models, verifiable flows, and explicit conformance requirements.
2.1  Definition
Agent Control Protocol (ACP) is a technical specification that defines the mechanisms by which autonomous institutional agents are identified, authorized, monitored, and governed in B2B environments. It establishes the formal contract between an agent, the institution that operates it, and the institutions with which it interacts.
Core invariant:
$\textit{Execute}(\text{request})\Longrightarrow\textit{ValidIdentity}(\text{agent})\;\wedge\;\textit{ValidCapability}\;\wedge\;\textit{ValidDelegationChain}\;\wedge\;\textit{AcceptableRisk}$
No action of an ACP agent can be executed without all four predicates being simultaneously true. If any fails, the action is denied. No exceptions.
2.2  Design Principles
ACP was designed with five principles that are non-negotiable at implementation time:
  • P1
    Fail Closed. On any internal component failure, the action is denied. Never approved by default.
  • P2
    Identity is cryptography. $\text{AgentID}=\text{base58}(\text{SHA-256}(\text{public\_key}))$. No usernames. No arbitrary IDs. Identity cannot be claimed—it must be demonstrated in every request.
  • P3
    Delegation does not expand privileges. The delegated agent’s permissions are always a strict subset of the delegator’s permissions. This property is cryptographically verified at every chain hop.
  • P4
    Complete auditability. Every decision—approved, denied, or escalated—is recorded in an append-only ledger with institutional signature. Not just successes. Everything.
  • P5
    External verification possible. Any institution can verify ACP artifacts from another institution using only the public key registered in the ITA. No dependency on proprietary systems.
2.3  Formal Agent Model
In ACP, an agent is a formal tuple with well-defined state:
$A=(\,\text{AgentID},\;\text{capabilities},\;\text{autonomy\_level},\;\text{state},\;\text{limits}\,)$
Field Type Description
AgentID String (43–44 chars) base58(SHA-256(pk)). Derived from public key. Immutable.
capabilities List of strings Explicit permissions. Format: acp:cap:<domain>.<action>. Never abstract roles.
autonomy_level Integer 0–4 Determines risk evaluation thresholds. 0 = no autonomy. 4 = maximum.
state Enum active | restricted | suspended | revoked. Transition to revoked is unidirectional.
limits Object Rate limits, maximum amounts, time windows. Not modifiable at runtime.
Table 2: ACP formal agent tuple fields.
2.4  Layered Architecture
ACP does not replace existing security infrastructure. It is added as an upper layer:
ACP Layer -- Autonomous agent governance: identity, authorization, risk, auditing
RBAC Layer -- Role-based access control for human users
Zero Trust -- Continuous identity and network access verification
3  Technical Mechanisms
ACP defines six interdependent mechanisms. Each has its own formal specification, state model, data structure, protocol flow, and error codes.
3.1  Serialization and Signing (ACP-SIGN-1.0)
Every verification in ACP begins with signature verification. ACP-SIGN-1.0 defines the exact process that produces a binary result—valid or invalid—without ambiguity:
  1. 1.
    Canonicalization with JCS (RFC 8785) [11]. Produces a deterministic representation of the JSON object, independent of field order and the system that generated it.
  2. 2.
    SHA-256 hash over the canonical output in UTF-8.
  3. 3.
    Ed25519 signature (RFC 8032) [7] over the hash. 32-byte key, 64-byte signature.
  4. 4.
    Base64url encoding without padding for transmission.
Signature verification precedes all semantic validation. An object with an invalid signature is rejected without processing its content. This rule has no exceptions (PROHIB-003, PROHIB-012).
3.2  Capability Token (ACP-CT-1.0)
The Capability Token is ACP’s central artifact. It is a signed JSON object that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
"ver": "1.0",
"iss": "<AgentID_issuer>",
"sub": "<AgentID_subject>",
"cap": ["acp:cap:financial.payment"],
"res": "org.example/accounts/ACC-001",
"exp": 1718923600,
"nonce": "<128bit_CSPRNG_base64url>",
"deleg": { "allowed": true, "max_depth": 2 },
"parent_hash": null,
"sig": "<Ed25519_base64url>"
Critical fields: exp is mandatory—a token without expiry is invalid by definition. The 128-bit nonce prevents replay attacks. parent_hash chains delegated tokens in a verifiable way. The signature covers all fields except sig.
3.3  Handshake and Proof-of-Possession (ACP-HP-1.0)
Possessing a valid Capability Token is not sufficient to act. ACP-HP-1.0 requires that the bearer demonstrate in every request that they possess the private key corresponding to the AgentID declared in the token. This eliminates the possibility of impersonating an agent with a stolen token.
The protocol is stateless—it does not establish sessions, does not produce a session_id, does not require server-side state between requests. The proof occurs in every interaction:
  1. 1.
    The receiving system issues a 128-bit challenge generated by CSPRNG, valid for 30 seconds and single-use.
  2. 2.
    The agent signs the challenge together with the HTTP method, path, and body hash of the request.
  3. 3.
    The receiver verifies the signature using the agent’s public key, obtained from the ITA.
  4. 4.
    The challenge is deleted immediately after use—it cannot be reused.
This sequence guarantees four formal properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
3.4  Deterministic Risk Evaluation (ACP-RISK-2.0)
Each authorization request passes through a deterministic risk function that produces a Risk Score (RS) in the range $[0,100]$. The same input always produces the same result—no stochastic elements, no machine learning in the critical path.
$\text{RS}=\min\!\bigl(100,\;B(c)+F_{\text{res}}(r)+F_{\text{ctx}}(x)+F_{\text{hist}}(h)+F_{\text{anom}}(q)\bigr)$
where $F_{\text{anom}}(q)$ is the anomaly factor introduced in v2.0, evaluated against the agent’s accumulated request history $q$ (pattern frequency, denial rate, cooldown state). When an agent accumulates three DENIED decisions within ten minutes, a cooldown is activated and subsequent requests are blocked regardless of RS.
Factor Description Example values
$B(c)$ Baseline by capability *.read = 0, financial.payment = 35
$F_{\text{ctx}}(x)$ Request context Non-corporate IP +20+20; outside hours +15+15
$F_{\text{hist}}(h)$ Agent history (24h) Recent denial +20+20; anomalous frequency +15+15
$F_{\text{res}}(r)$ Resource classification public = 0; sensitive = 15; restricted = 45
Table 3: ACP-RISK-2.0 risk scoring factors (base formula; $F_{\text{anom}}$ adds up to +30+30 from pattern accumulation).
The RS determines the decision according to the thresholds configured for the agent’s autonomy_level. With autonomy_level 2 (standard): $\text{RS}\leq 39$ $\rightarrow$ APPROVED; $\text{RS}\in[40,69]$ $\rightarrow$ ESCALATED; $\text{RS}\geq 70$ $\rightarrow$ DENIED. An agent with autonomy_level 0 always receives DENIED.
3.5  Verifiable Chained Delegation (ACP-DCMA-1.1)
ACP allows an agent to delegate capabilities to another agent, which in turn can delegate to a third, up to the maximum depth defined in the root token. Delegation guarantees three properties:
  • No privilege escalation. The delegated agent’s capability set is always $\subseteq$ the delegator’s set. Cryptographically verified at every hop via the parent_hash field.
  • Bounded depth. The max_depth field establishes the chain limit. A chain that exceeds it is invalid.
  • Transitive revocation. Revoking an agent’s token automatically invalidates all delegated tokens that descend from it.
3.6  Execution Token (ACP-EXEC-1.0)
The separation between authorization and execution is a core principle of ACP. When the authorization engine approves a request, it returns an Execution Token (ET): a single-use artifact with a short lifetime that authorizes exactly that action, on that resource, at that moment.
  • An ET can only be consumed once. A second presentation is rejected (PROHIB-002).
  • An expired ET is invalid even if it was never used.
  • The target system that consumes the ET notifies the ACP endpoint, closing the audit cycle.
3.7  Audit Ledger (ACP-LEDGER-1.3)
The Audit Ledger is a chain of cryptographically signed events where each event includes the hash of the previous event:
$h_{n}=\text{SHA-256}(\,e_{n}\;\|\;h_{n-1}\,)$
This structure makes it impossible to modify or delete an event without invalidating the entire subsequent chain. The ledger records all lifecycle event types: GENESIS, AUTHORIZATION (including DENIED and ESCALATED), RISK_EVALUATION, TOKEN_ISSUED, TOKEN_REVOKED, EXECUTION_TOKEN_ISSUED, and EXECUTION_TOKEN_CONSUMED.
Institutions with FULL conformance expose the ledger via GET /acp/v1/audit/query, allowing external partners to verify chain integrity using only the institutional ITA public key. Modifying or deleting ledger events is prohibited (PROHIB-007, PROHIB-008).
It is important to note that the ledger provides verifiable evidence of execution, not enforcement itself. Enforcement in ACP occurs at the policy evaluation and execution layers (ACP-EXEC-1.0, ACP-DCMA-1.1): the Execution Token is the cryptographic artifact that gates actual system-state mutation. The ledger’s role is to make every admission decision—and its outcome—auditable and tamper-evident after the fact.
4  Inter-Institutional Trust
In a B2B environment, agents of one institution interact with another institution’s systems. ACP defines the exact mechanism by which this trust is established, verified, and can be revoked.
4.1  Institutional Trust Anchor (ACP-ITA-1.0)
The ITA is the authoritative registry that links an institution_id to an Ed25519 public key. It is the only point where ACP depends on an out-of-band mechanism: the initial distribution of the ITA authority’s public key. Once that key is resolved, all subsequent verification is autonomous and cryptographic.
Each institution registers in the ITA its Root Institutional Key (RIK)—the private key held in HSM that never leaves it. All ACP artifacts from that institution (tokens, ledger events, API responses) are signed with that key. Any third party can verify them by resolving the public key from the ITA.
4.2  Mutual Recognition (ACP-ITA-1.1)
When two institutions operate under different ITA authorities, ACP-ITA-1.1 defines the mutual recognition protocol. The process requires both authorities to sign a Mutual Recognition Agreement (MRA) establishing: the scope of recognition, the agreement’s validity period, and the proxy resolution mechanism.
Recognition is explicitly non-transitive. If A recognizes B and B recognizes C, A does not automatically recognize C. Each bilateral relationship requires its own signed MRA. This prevents uncontrolled expansion of the trust graph.
4.3  Institutional Key Rotation and Revocation
Normal rotation includes a transition period of up to 7 days during which both keys are valid, allowing artifacts signed with the old key to be verified during the transition.
Emergency rotation is activated when a key is compromised. The result is immediate: the key is marked revoked, all artifacts signed with it are invalid from that moment, and there is no transition period.
5  Cross-Organization Execution
ACP-CROSS-ORG-1.1 defines a fault-tolerant bilateral protocol for interactions between independent institutions, each operating its own ACP ledger. The protocol closes five gaps identified in the 1.0 version: undefined status tracking, absent retry protocol, unregistered ACK event type, pending_review without SLA, and a stale header reference.
5.1  Interaction Model
A cross-organization interaction involves two ACP-compliant institutions: a source institution that originates the bundle and a target institution that validates, executes, and acknowledges it. The protocol operates asynchronously—the source does not block waiting for an ACK.
Each interaction carries a mandatory interaction_id (UUIDv7), immutable and reused across all retries. This is distinct from event_id (UUIDv4), which is unique per emission. The target deduplicates by interaction_id, not by event_id.
For example, an agent in Institution A may request execution in Institution B. The request is recorded as a CROSS_ORG_INTERACTION event in A’s ledger, transmitted to B, validated against B’s policy and capability registry, and resolved through a CROSS_ORG_ACK event registered in both ledgers. Institution A derives final execution state from the ACK without storing mutable status.
5.2  Fault Tolerance and Retry Protocol
When no CROSS_ORG_ACK is received within the 300-second timeout, the source retries up to three times with exponential backoff:
Attempt Wait after timeout
1 (initial)
2 +30 s
3 +60 s
4 (final) +120 s
After three failed attempts the interaction transitions to retry_exhausted (CROSS-012) and an operational alert is raised. If a valid ACK arrives at any point, all retry timers are cancelled immediately.
5.3  Derived Interaction Status
The status of a cross-organization interaction is never stored as mutable state. It is always derived from events present in the ledger:
Derived status Condition
pending_ack CROSS_ORG_INTERACTION exists, no ACK
acked ACK with status = accepted
rejected ACK with status = rejected
pending_review ACK with status = pending_review
expired pending_review AND now > review_deadline
Precedence rule: accepted >> rejected >> pending_review. A mutable CROSS_ORG_STATUS_UPDATE event is explicitly prohibited as an anti-pattern that introduces race conditions and divergent audit trails.
5.4  Interaction State Invariants
For any interaction_id, ACP enforces the following invariants:
  • At most one terminal state. Only one accepted or rejected ACK may exist per interaction_id; a duplicate with a different payload is rejected (CROSS-014).
  • ACK dominates retries. A valid CROSS_ORG_ACK cancels all pending retry timers immediately, regardless of attempt number.
  • Immutable correlation key. Retry operations MUST NOT create new interaction_id values; only event_id changes across attempts.
  • No terminal-to-non-terminal regression. Once accepted or rejected, the status cannot change. The transition pending_review $\to$ pending_review is also prohibited (CROSS-015).
5.5  Pending Review SLA
When the target returns pending_review, a 24-hour SLA applies. The ACK payload includes a review_deadline field (Unix seconds). Valid terminal transitions are: accepted, rejected, or expired (implicit, when now > review_deadline).
5.6  CROSS_ORG_ACK as a First-Class Ledger Event
CROSS_ORG_ACK is registered in ACP-LEDGER-1.3 §5.15 as a first-class event type, signed with Ed25519 and serialized via JCS (RFC 8785). This enables verifiable and auditable execution across independent institutional boundaries, eliminating the audit gap that existed in version 1.0.
5.7  Security Considerations
ACP-CROSS-ORG-1.1 assumes authenticated communication channels between institutions. All CROSS_ORG_INTERACTION and CROSS_ORG_ACK events are signed with the originating institution’s Ed25519 ITA key and serialized via JCS (RFC 8785), providing authenticity, integrity, and non-repudiation.
Without this cryptographic layer, cross-organization events would be subject to forgery, replay, and unauthorized status derivation. Specifically, ACP mitigates:
  • Replay attacks: interaction_id (UUIDv7) and event_id (UUIDv4) allow the target to detect and reject duplicate submissions.
  • Event forgery: Ed25519 signatures over JCS-canonical payloads prevent an adversary from injecting valid-looking ACK or interaction events.
  • Audit divergence: Because CROSS_ORG_ACK is a first-class ledger event on both sides, independent auditors can verify consistency without trusting either institution’s mutable state.
ACP does not eliminate all sources of failure in distributed systems, but provides a structured model for detecting, handling, and auditing them.
6  Reputation Snapshot Portability (ACP-REP-PORTABILITY-1.1)
ACP-REP-PORTABILITY-1.1 defines the ReputationSnapshot: a compact, cryptographically signed record that carries an agent’s reputation score across organizational boundaries. Unlike the bilateral attestation protocol of v1.0, this specification focuses on the snapshot object itself—its structure, signing procedure, validation algorithm, and expiration semantics—enabling any verifier to independently validate a snapshot without trusting an intermediary.
6.1  ReputationSnapshot Structure
A ReputationSnapshot contains ten fields: ver (version), rep_id (UUID v4), subject_id, issuer, score (float), scale ("0-1" or "0-100"), model_id, evaluated_at (Unix seconds), valid_until (Unix seconds), and signature (Ed25519, base64url). Fields scale, model_id, and valid_until are new in v1.1; v1.0 snapshots omit them and are accepted without expiration enforcement (backward compatibility, §12).
6.2  Signing Procedure
The signature covers all fields except signature itself, serialized via JCS (RFC 8785) canonicalization, SHA-256 digest, and Ed25519 signing:
  1. 1.
    raw = json.Marshal(signableReputation)
  2. 2.
    canonical = jcs.Transform(raw) // RFC 8785
  3. 3.
    digest = SHA-256(canonical)
  4. 4.
    sig = Ed25519.Sign(privKey, digest)
  5. 5.
    snapshot.signature = base64url(sig) // no padding
JCS canonicalization is mandatory. Using json.Marshal directly is prohibited: key ordering differs across Go, Python, and TypeScript implementations, producing verification failures in cross-org deployments.
6.3  Validation Invariants
The validator enforces five invariants (REP-001 through REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
Structural validation (Validate) and cryptographic verification (VerifySig) are intentionally separate operations, enabling lightweight ingestion-time checks without requiring the issuer’s public key.
6.4  Divergence Semantics
When a verifier receives snapshots of the same subject_id from multiple issuers, it may compute divergence: $|a.\text{score}-b.\text{score}|$. If the divergence exceeds a configurable threshold (default 0.30 for scale="0-1"), the verifier emits warning REP-WARN-002. This is non-blocking—the policy decision of whether to accept, escalate, or reject remains with the verifier’s business logic, preserving institutional sovereignty.
7  Multi-Organization Interoperability Demo (GAP-14)
The examples/multi-org-demo/ directory provides an executable end-to-end demonstration of cross-organizational ACP exchange, combining ACP-POLICY-CTX-1.1 and ACP-REP-PORTABILITY-1.1 in a single runnable scenario deployable with Docker in under five minutes.
7.1  Scenario
Two independent organizations, Org-A and Org-B, interact over HTTP. Org-A is the issuer: it evaluates an agent request against its local policy and produces two signed artifacts—a PolicyContextSnapshot and a ReputationSnapshot—using its Ed25519 institutional key. Org-B is the verifier: it independently validates both artifacts using the pkg/policyctx and pkg/reputation packages (no logic duplicated) and renders an autonomous admission decision.
The validation sequence in Org-B is:
  1. 1.
    policyctx.VerifySig(pcs, pubKey) — verifies institutional signature (ACP-POLICY-CTX-1.1 §6 step 12).
  2. 2.
    policyctx.VerifyCaptureFreshness(pcs, 300s) — enforces $\delta_{\max}$ freshness invariant.
  3. 3.
    reputation.Validate(rep, now) — structural invariants REP-001 through REP-011.
  4. 4.
    reputation.VerifySig(rep, pubKey) — Ed25519/JCS/SHA-256 cryptographic verification.
  5. 5.
    reputation.CheckDivergence(orgA, orgB, 0.30) — divergence check against Org-B’s own score.
  6. 6.
    Final ACCEPT / DENY decision under Org-B’s sovereign policy.
7.2  Divergence Handling and Institutional Sovereignty
If Org-B holds its own score for the same subject_id, it computes divergence $|a.\text{score}-b.\text{score}|$. If this exceeds 0.30, Org-B emits REP-WARN-002. REP-WARN-002 is non-blocking—the final decision is Org-B’s prerogative, independent of Org-A’s assessment.
This embodies the ACP invariant: ACP reports divergence. ACP does not resolve divergence. Org-B cannot extend Org-A’s valid_until, override Org-A’s delta_max, or substitute its own snapshot for Org-A’s signed record. Each institution’s attestations are cryptographically bound to its own key and immutable after issuance.
7.3  Deployment
The demo ships with a single Dockerfile (parameterized by ARG ORG=org-a|org-b), a docker-compose.yml with health check, and a dedicated Go module with a replace directive pointing to impl/go/:
docker compose up --build # from examples/multi-org-demo/
curl http://localhost:8081/request | jq .
8  Security Model
ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope.
8.1  Threat Model (STRIDE)
Category Threat Mitigation in ACP
Spoofing AgentID impersonation $\text{AgentID}=\text{SHA-256}(\text{pk})$. Without valid signature $\rightarrow$ immediate DENIED.
Tampering Token or event alteration Ed25519 covers all fields. Chained ledger: altering one event invalidates all subsequent.
Repudiation Agent denies executed action ActionRequest digitally signed. Non-repudiation by design.
Info Disclosure Capability exposure Tokens reveal only the necessary subset. Channel confidentiality via TLS.
DoS Request flooding Rate limits per agent_id. Anomalous frequency control.
Elevation Delegation that expands privileges $C_{\text{delegated}}\subseteq C_{\text{original}}$. Cryptographically verified at each hop.
Table 4: STRIDE threat model and ACP mitigations.
8.2  Guaranteed Security Properties
ACP guarantees the following properties when the implementation is compliant:
  • Artifact integrity. EUF-CMA security of Ed25519. Impossible to modify a token or event without invalidating the signature.
  • Identity authenticity. Only whoever possesses sk can generate a valid signature under the corresponding pk.
  • No privilege escalation via delegation. Demonstrable by induction over the delegation chain.
  • Anti-replay. The single-use challenge makes reusing a proof of possession ineffective.
  • Effective revocation. $\textit{Valid}(t)=\textit{valid\_sig}\wedge\neg\textit{expired}\wedge\neg\textit{revoked}\wedge\textit{valid\_delegation}$. All four conditions must be true simultaneously.
8.3  Declared Residual Risks
ACP explicitly declares what it cannot resolve:
  • Total compromise of the RIK held in HSM. ACP defines the emergency rotation process but cannot prevent a physical compromise of the custody infrastructure.
  • Coordinated institutional collusion. If multiple institutions act maliciously in a coordinated manner, they can generate valid artifacts. ACP guarantees traceability, not prevention of malicious agreements between parties.
  • Implementation failures. ACP is a specification. An implementation that violates prohibited behaviors can compromise all protocol guarantees. Conformance requires formal testing.
  • ITA bootstrap. The only point that depends on an out-of-band channel. Once the ITA root key is resolved, everything subsequent is autonomous.
9  Conformance and Interoperability
9.1  Conformance Levels
Level Required documents Enabled capability
L1 — CORE ACP-SIGN-1.0, ACP-CT-1.0, ACP-HP-1.0 Token issuance with cryptographic PoP
L2 — SECURITY L1 + ACP-RISK-2.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (F_anom + Cooldown), transitive revocation, ITA
L3 — FULL L2 + ACP-API-1.0, ACP-EXEC-1.0, ACP-LEDGER-1.3 Complete verifiable auditing
L4 — EXTENDED L3 + ACP-PAY-1.0, ACP-NOTIFY-1.0, ACP-DISC-1.0, … Full governance suite
L5 — DECENTRAL. L4 + ACP-D suite Federation without central ITA
Table 5: ACP-CONF-1.2 conformance levels.
9.2  Conformance Declaration
Every compliant implementation MUST expose a public endpoint without authentication:
GET https://<endpoint>/acp/v1/conformance
This endpoint returns the achieved level, implemented documents, declared extensions, and declaration date. It allows any external partner to verify the conformance level of a counterparty before establishing an ACP relationship.
9.3  Prohibited Behaviors
ACP defines 12 behaviors that no compliant implementation can exhibit. If any is exhibited, the implementation cannot declare conformance at any level:
Code Prohibited behavior
PROHIB-001 Approving a request when any evaluation component fails
PROHIB-002 Reusing an already-consumed Execution Token
PROHIB-003 Omitting signature verification on any incoming artifact
PROHIB-004 Treating a not-found token_id as active in a revocation context
PROHIB-005 Allowing state transition from revoked
PROHIB-006 Issuing an ET without a prior APPROVED AuthorizationDecision
PROHIB-007 Modifying or deleting Audit Ledger events
PROHIB-008 Silencing ledger corruption detection
PROHIB-009 Ignoring max_depth in delegation chains
PROHIB-010 Implementing an offline policy more permissive than ACP-REV-1.0
PROHIB-011 Approving requests from agents with autonomy_level 0
PROHIB-012 Continuing to process an artifact with an invalid signature
Table 6: ACP prohibited behaviors. Exhibiting any disqualifies conformance at all levels.
9.4  B2B Interoperability Conditions
  • L1 Interoperability: Institution A can verify tokens from B if both implement ACP-CONF-L1, A has access to B’s public key (via ITA or out-of-band), and B’s tokens use ACP-SIGN-1.0 algorithms.
  • L2 Interoperability: A can delegate to B’s agents if both implement ACP-CONF-L2, are registered in a common ITA or with mutual recognition, and B’s revocation endpoint is accessible to A.
  • L3 Interoperability: A can audit B’s ledger if B implements ACP-CONF-L3, A can resolve B’s public key via ITA, and B exposes GET /acp/v1/audit/query.
10  Use Cases
ACP is sector-agnostic. The mechanisms are identical regardless of industry. What varies is the configuration of capabilities, resources, and autonomy levels.
10.1  Financial Sector — Inter-Institutional Payment Agents
ACP-PAY-1.0 extends the capability registry with formal specifications for acp:cap:financial.payment and acp:cap:financial.transfer. Mandatory constraints in the token include max_amount and currency. 12 specific validation steps cover limit verification, beneficiary validation, and time window control. In a financial B2B scenario, Bank A can authorize an agent to execute payments up to a defined amount to pre-approved beneficiaries, with a complete record verifiable by Bank B without needing shared proprietary systems.
10.2  Digital Government — Document Processing
Government agents that process documents can operate under ACP with autonomy_level 1 or 2, requiring human review for any action with a Risk Score above the configured threshold. Ledger traceability provides forensic evidence for regulatory audits and transparency processes.
10.3  Enterprise AI — Multi-Company Orchestration
In agent pipelines that cross organizational boundaries, ACP allows each organization to maintain formal control over what other organizations’ agents can do in their systems. Chained delegation allows an agent in company A to operate in company B’s systems with explicitly delegated capabilities, without B needing to trust A’s internal controls—only the chain of signed tokens.
10.4  Critical Infrastructure — Monitoring and Actuation Agents
For systems where an incorrect action has irreversible consequences, autonomy_level 0 ensures any actuation request is DENIED without evaluating the Risk Score. The ledger provides the forensic record needed for post-incident analysis.
11  Specification and Implementation Status
ACP v1.17 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
11.1  Active Specifications — v1.17 (38 documents)
Level Document Title
L1 ACP-SIGN-1.0 Serialization and Signature
L1 ACP-AGENT-1.0 Agent Identity
L1 ACP-CT-1.0 Capability Tokens
L1 ACP-CAP-REG-1.0 Capability Registry
L1 ACP-HP-1.0 Handshake / Proof-of-Possession
L1 ACP-DCMA-1.1 Delegated Chain Multi-Agent
L1 ACP-MESSAGES-1.0 Wire Message Format
L1 ACP-PROVENANCE-1.0 Authority Provenance
L1 ACP-SIGN-2.0 Hybrid Signature (Ed25519 + ML-DSA-65)
L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (F_anom + Cooldown)
L2 ACP-RISK-1.0 Deterministic Risk Engine v1 (superseded by v2)
L2 ACP-REV-1.0 Revocation Protocol
L2 ACP-ITA-1.0 Institutional Trust Anchor
L2 ACP-ITA-1.1 ITA Mutual Recognition
L2 ACP-REP-1.2 Reputation Module
L3 ACP-API-1.0 HTTP API
L3 ACP-EXEC-1.0 Execution Tokens
L3 ACP-LEDGER-1.3 Audit Ledger (mandatory institutional sig)
L3 ACP-PSN-1.0 Policy Snapshot
L3 ACP-POLICY-CTX-1.1 Policy Context Snapshot
L3 ACP-LIA-1.0 Liability Attribution
L4 ACP-HIST-1.0 History Query API
L4 ACP-PAY-1.0 Financial Capability
L4 ACP-NOTIFY-1.0 Event Notifications
L4 ACP-DISC-1.0 Service Discovery
L4 ACP-BULK-1.0 Batch Operations
L4 ACP-CROSS-ORG-1.1 Cross-Organization Bundles
L4 ACP-GOV-EVENTS-1.0 Governance Event Stream
L4 ACP-REP-PORTABILITY-1.1 Reputation Snapshot Portability
Gov. ACP-CONF-1.2 Conformance — sole normative source
Gov. ACP-TS-1.1 Test Vector Format
Gov. RFC-PROCESS Specification Process
Gov. RFC-REGISTRY Specification Registry
Gov. ACP-CR-1.0 Change Request Process
Table 7: ACP v1.17 active specification documents by conformance level (38 primary documents shown). The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published. New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec). New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/). Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
11.2  Reference Implementation — 23 Go Packages + Demos
The Go reference implementation in impl/go/ covers all L1–L4 conformance levels. All 23 packages pass go test ./.... A Python SDK (impl/python/) covers the ACP-HP-1.0 handshake and all ACP-API-1.0 endpoints. The examples/multi-org-demo/ directory (GAP-14) provides a runnable two-organization scenario using the real pkg/policyctx and pkg/reputation packages; see Section 7. The examples/payment-agent/ directory (Sprint E) provides an executable HTTP server demonstrating the full ACP-RISK-2.0 pipeline: POST /admission evaluates the request with the deterministic risk engine, appends an immutable ledger event with the full factor breakdown, and triggers cooldown automatically after three DENIED decisions in ten minutes.
Package Spec Level
pkg/crypto ACP-SIGN-1.0 + ACP-AGENT-1.0 L1
pkg/tokens ACP-CT-1.0 L1
pkg/registry ACP-CAP-REG-1.0 L1
pkg/handshake ACP-HP-1.0 L1
pkg/delegation ACP-DCMA-1.1 L1
pkg/provenance ACP-PROVENANCE-1.0 L1
pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 L2
pkg/revocation ACP-REV-1.0 L2
pkg/reputation ACP-REP-1.2 + ACP-REP-PORTABILITY-1.1 L2/L4
pkg/execution ACP-EXEC-1.0 L3
pkg/ledger ACP-LEDGER-1.3 L3
pkg/psn ACP-PSN-1.0 L3
pkg/policyctx ACP-POLICY-CTX-1.1 L3
pkg/lia ACP-LIA-1.0 L3
pkg/hist ACP-HIST-1.0 L4
pkg/govevents ACP-GOV-EVENTS-1.0 L4
pkg/notify ACP-NOTIFY-1.0 L4
pkg/disc ACP-DISC-1.0 L4
pkg/bulk ACP-BULK-1.0 L4
pkg/crossorg ACP-CROSS-ORG-1.1 L4
pkg/pay ACP-PAY-1.0 L4
cmd/acp-server ACP-API-1.0 L3
Table 8: Go reference implementation packages (23 total). All pass go test ./....
11.3  Conformance Test Vectors — 73 Signed Vectors
The compliance/test-vectors/ directory contains 73 signed test vectors per ACP-TS-1.1:
Suite Positive Negative Spec
CORE (SIGN, CT, HP) 4 4 L1
DCMA 2 2 L1
HP 2 8 L1
LEDGER 3 8 L3
EXEC 2 7 L3
PROV 2 7 L1
PCTX 4 9 L3
REP 3 6 L4
Total 22 51
Table 9: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A) and real SHA-256 hash chains.
11.4  ACP-RISK-2.0 Engine Performance
The pkg/risk ACP-RISK-2.0 engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8). These measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
Benchmark ns/op B/op allocs/op
Evaluate (APPROVED, no anomaly) 1,012 280 5
Evaluate (DENIED, no anomaly) 863 296 5
Evaluate (all 3 F_anom rules) 1,331 296 5
Evaluate (COOLDOWN short-circuit) 149 112 1
PatternKey (SHA-256 hash) 996 184 4
ShouldEnterCooldown 88 0 0
[…]
The deterministic design of ACP-RISK-2.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable latency bounds. At these throughput characteristics, the admission check adds negligible overhead relative to the actions it governs.
12  Evaluation
This section evaluates ACP along three dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, and (iii) reproducibility and external verifiability of decision outcomes.
12.1  Evaluation Goals
  • Q1: What is the computational overhead of ACP decision evaluation relative to the actions it governs?
  • Q2: Does the system produce correct decisions under stateful conditions such as repeated denials, pattern accumulation, and cooldown activation?
  • Q3: Can ACP executions be deterministically reproduced and externally verified without access to internal implementation details?
12.2  Experimental Setup
All experiments use the Go reference implementation (v1.17) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8). The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13). Production deployments with persistent storage backends will incur additional I/O latency not measured here.
12.3  Q1: Computational Overhead
Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 1,331 ns. The cooldown short-circuit path (149 ns) bypasses the risk function entirely.
Answer to Q1: The admission check adds between 149 ns and 1,331 ns of latency per request. Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
12.4  Q2: Stateful Behavior Correctness
The ACR-1.0 compliance runner executes each sequence vector end-to-end under the execution contract of Section 13. Results:
Vector Behavior tested Result
SEQ-BENIGN-001 No false positives under repeated reads PASS
SEQ-BOUNDARY-001 Exact thresholds RS = 0, 25, 35, 40, 70 PASS
SEQ-PRIVJUMP-001 Low→high privilege jump caught immediately PASS
SEQ-FANOM-RULE3-001 Pattern count $\geq 3$ adds +15+15 at step 4 PASS
SEQ-COOLDOWN-001 3 DENIED in 10 min → cooldown active PASS
Total 5/5 PASS
Table 11: ACR-1.0 sequence vector results under strict mode (--strict).
Answer to Q2: All five stateful scenarios pass with CONFORMANT status. State transitions follow the defined execution contract; cooldown activates deterministically after the threshold is crossed; pattern accumulation influences decisions exactly as specified in ACP-RISK-2.0 Rule 3.
12.5  Q3: Reproducibility and External Verifiability
Given the same sequence of inputs and the same initial state, the ACR-1.0 runner produces identical outputs on every run. This holds because:
  • Evaluate() is a pure function of its arguments and the querier state,
  • state transitions follow the deterministic execution contract,
  • and the test vectors ship with the specification, enabling any third party to reproduce results without access to proprietary infrastructure.
Answer to Q3: ACP executions are fully reproducible and externally verifiable. This is a property not typically supported by conventional policy engines, which evaluate policy against a snapshot of state without defining how that state evolves across requests [12, 6, 9].
12.6  Limitations
The evaluation uses an in-memory state backend. Persistent storage, distributed deployments, and concurrent request patterns introduce latency and consistency considerations not measured here. Adversarial robustness under coordinated attack scenarios is not evaluated. These are identified as future work (Section 12 roadmap).
12.7  Roadmap
Item Status
Core specs (L1–L4), 38 documents Complete
Go reference implementation (23 packages) Complete
Conformance test vectors (73 signed + 65 unsigned RISK-2.0) Complete
OpenAPI 3.1.0 (openapi/acp-api-1.0.yaml, 18 endpoints) Complete
Python SDK (impl/python/) Complete (L1 + full API client)
Docker image (ghcr.io/chelof100/acp-server:latest) Complete
Multi-org interoperability demo (examples/multi-org-demo/) Complete (GAP-14)
ACP-RISK-2.0 (F_anom + Cooldown, pkg/risk) v1.16 Complete
Payment-agent demo (examples/payment-agent/) v1.16 Complete
ACP-SIGN-2.0 spec (Ed25519 + ML-DSA-65 hybrid) v1.16 Complete
ACR-1.0 sequence compliance runner + 5 stateful vectors v1.17 Complete
TLA+ formal model (TLC-runnable, 3 invariants) v1.17 Complete
ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 Complete
Post-quantum Go implementation (Dilithium, circl) v1.18
L5 Decentralized (ACP-D) Specification in design (v2.0)
IETF RFC submission After L5 stabilization
Table 12: ACP v1.17 development roadmap.
13  Compliance Testing (ACR-1.0 Sequence Runner)
We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 implementations against sequence-based test cases. The runner operates in two modes: library mode (direct call to pkg/risk, no network, deterministic) and HTTP mode (external server validation for interoperability). All 73 signed single-shot conformance vectors plus 5 sequence scenarios are validated on every commit.
13.1  Sequence-Based Test Vectors
Unlike single-shot conformance vectors, sequence-based vectors capture temporal dynamics and emergent risk patterns across multiple requests to the same engine instance. The compliance runner executes steps in order, maintains state between them (request history, pattern counts, denial counts, cooldown), and validates each step’s decision and risk_score against expected values.
The five sequence vectors in compliance/test-vectors/sequence/ cover:
Vector ID Steps Behavior Tested Key Invariant
SEQ-BENIGN-001 3 Benign repeated reads Baseline: RS=0, no state buildup
SEQ-BOUNDARY-001 3 RS boundary conditions Exact thresholds: 0, 25, 35, 40, 70
SEQ-PRIVJUMP-001 2 Low-risk → high-risk jump No residual benefit from prior APPROVED
SEQ-FANOM-RULE3-001 4 F_anom Rule 3 activation Pattern count $\geq 3$ at step 4 adds +15+15
SEQ-COOLDOWN-001 4 Cooldown trigger and block 3 DENIED in 10 min → cooldown active
Table 13: ACR-1.0 sequence test vectors. All 5/5 pass in library mode with the Go reference implementation.
Execution contract. The runner implements the execution contract defined by ACP-RISK-2.0: evaluate first (stateless), then update state. Critically, AddPattern() is called after Evaluate(), so the pattern count visible to step $n$’s evaluation is n1n-1. F_anom Rule 3 (pattern count $\geq 3$) therefore first triggers on step 4, not step 3. This contract is what makes the system testable and verifiable by third parties.
now := time.Now() // capture once
result, _ := risk.Evaluate(req, querier) // 1: stateless eval
querier.AddRequest(req.AgentID, now) // 2: always
querier.AddPattern(patKey, now) // 3: always (feeds F_anom Rule 3)
if result.Decision == risk.DENIED {
querier.AddDenial(req.AgentID, now) // 4: conditional
if should, _ := risk.ShouldEnterCooldown(...) // 5: check threshold
querier.SetCooldown(agentID, now.Add(p)) // 6: set expiry
Listing 1: ACR-1.0 execution contract in library mode.
13.2  Model–Implementation Alignment
The TLA+ model, test vectors, and runtime evaluation are aligned through a shared request abstraction:
TLA+ variable Runner field Engine field
capability, resource RunnerRequest EvalRequest
risk_score risk_score RSFinal
decision decision Decision
ledger (audit log) InMemoryQuerier events
cooldown denied_reason CooldownActive()
Table 14: Alignment across TLA+ formal model, ACR-1.0 runner, and ACP-RISK-2.0 engine.
14  End-to-End Verifiability
ACP v1.17 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction. Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
TLA+ ModelSafety $\wedge$ AppendOnly $\wedge$ DeterminismTest Vectorssequence/ + single-shotComplianceRunner (ACR-1.0)ACP Enginepkg/riskAudit Ledgerappend-onlydefinesinputcallsrecords statically checked by TLC (0 violations, bounded state space)
Figure 1: ACP end-to-end verifiability pipeline. The TLA+ model defines the invariants that test vectors instantiate and the compliance runner validates at runtime. The ACP engine records every decision in the append-only audit ledger. The TLA+ layer is a static verification artifact; it does not run at admission time.
What is formally verified in TLA+ holds at runtime. Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner. This is the connection across the three layers of the v1.17 verifiability stack: formal $\to$ data $\to$ runtime.
15  How to Implement ACP
ACP is a specification—it does not require adopting any specific platform. It can be implemented on top of existing infrastructure.
15.1  Minimum Requirements for L1 Conformance
  • Ed25519 public key infrastructure. Key pair for each agent.
  • JCS implementation (RFC 8785). Deterministic canonicalization for all signed artifacts.
  • Capability Token issuance and verification with all mandatory fields per ACP-CT-1.0 §5.
  • Handshake endpoint to issue and verify challenges (ACP-HP-1.0 §6).
  • Capability registry with the core domains of ACP-CAP-REG-1.0.
15.2  Additional Requirements for L3 Conformance
  • Root Institutional Key held in HSM with documented rotation process.
  • Registration with an ITA authority (centralized or federated model).
  • Deterministic risk engine implementing ACP-RISK-2.0 (F_res, F_ctx, F_hist, F_anom, Cooldown).
  • Revocation endpoint (Mechanism A: online endpoint, or Mechanism B: CRL).
  • Append-only storage for the Audit Ledger with per-event signing.
  • Complete HTTP API per ACP-API-1.0, including health and conformance endpoints.
  • Public conformance declaration at GET /acp/v1/conformance.
15.3  What ACP Does Not Prescribe
ACP defines the what—mechanisms, flows, data structures, requirements. It does not prescribe the how of internal implementation: programming language, database, HSM provider, ITA provider, or specific integration with existing RBAC or Zero Trust systems.
16  Related Work
16.1  Access Control and Policy Enforcement
Traditional access control models such as Role-Based Access Control (RBAC) [12] and Attribute-Based Access Control (ABAC) [6] provide structured mechanisms for regulating access decisions based on predefined roles or attributes. Policy languages such as XACML [9] and modern policy engines such as Open Policy Agent [10] and Cedar [1] extend these paradigms with expressive evaluation engines.
These systems are fundamentally designed around stateless decision evaluation: they evaluate a request against policy at a single point in time. They do not define how system state evolves as a result of decisions over time, and they do not formalize how historical behavior (repeated denials, temporal patterns) influences subsequent decisions in a deterministic and auditable manner.
ACP differs by explicitly separating: (i) stateless decision evaluation, (ii) externalized state management, and (iii) an execution contract governing state evolution across requests. This enables reasoning not only about individual decisions, but about the temporal behavior of the admission system.
16.2  Runtime Policy Systems and Agent Frameworks
Recent protocols for autonomous agents and multi-agent coordination define structured interactions between components and services. The Model Context Protocol (MCP) [2] provides structured tool access between LLM applications and services. Agent-to-Agent (A2A) [4] defines communication and task delegation between agents.
While these protocols address interoperability and orchestration, they do not define:
  • a deterministic risk evaluation model that accumulates state across requests,
  • explicit state evolution semantics with a testable execution contract, or
  • externally verifiable execution traces that can be reproduced by third parties.
ACP complements these approaches by providing a governance layer that is independent of orchestration protocols, focusing specifically on verifiable decision enforcement and auditability.
16.3  Auditability and Tamper-Evident Logs
Secure audit logging [13] and event sourcing [3] ensure that past events can be reconstructed and verified after the fact. These approaches operate post hoc: they provide evidence of what occurred without constraining system behavior in real time.
ACP integrates auditability directly into the admission process. Every decision—APPROVED, ESCALATED, or DENIED—is recorded in an append-only, cryptographically chained ledger (ACP-LEDGER-1.3) before the governed action executes. This enables both real-time enforcement and post-hoc verification from the same artifact. The LedgerAppendOnlyTemporal invariant in the TLA+ model formally specifies that ledger entries are never modified or removed.
16.4  Formal Verification of Protocols
Model checking [8] and symbolic reasoning are used to establish correctness properties in distributed systems and protocols, typically focusing on safety and liveness of state transitions.
ACP incorporates formal methods in a targeted manner: the TLA+ model (tla/ACP.tla) mechanically checks three critical invariants (fail-closed admission, append-only ledger, risk determinism) using TLC. Rather than attempting full-system verification, ACP adopts a pragmatic approach in which formally verified invariants are instantiated as concrete sequence test vectors and validated at runtime by the ACR-1.0 compliance runner. This creates a three-layer verifiability chain: formal $\to$ data $\to$ runtime.
16.5  Summary
Existing systems address individual aspects of the problem—policy evaluation, agent orchestration, audit logging, or formal verification. ACP’s contribution lies in combining these elements into a unified model centered on an explicit execution contract. This enables a property not directly provided by prior systems: the ability to deterministically reproduce and externally verify the behavior of an admission decision system over time.
17  Conclusion
Autonomous agents are already operating in institutional environments. The question is not whether they will operate—it is whether they will do so with or without formal governance. ACP proposes that they operate with formal governance, with verifiable mechanisms, and with traceability that can withstand an external audit.
ACP is not the first attempt to control autonomous agents. It is the first attempt to do so through a formal technical specification with precise state models, demonstrable security properties, and verifiable conformance requirements. The difference between a best-practices policy and a formal protocol is exactly that: behaviors are defined, failures have specific error codes, and conformance can be verified.
The goal of ACP is not to make agents more capable. It is to make them governable. That is a necessary condition for their institutional deployment to be sustainable at scale.
ACP does not eliminate all sources of failure in distributed agent systems. It does not prescribe network transport, consensus mechanisms, or business-level dispute resolution. What it provides is a structured, auditable, and formally specified model for admission control, delegation, revocation, and cross-organization interaction— the layer that existing identity and policy frameworks leave unaddressed.
The v1.17 specification is complete. A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19198298
The specification and implementation are open for technical review, pilot implementation, and formal standardization.
TraslaIA invites organizations interested in adopting ACP, contributing to its evolution, or participating in the standardization process to reach out directly.
Marcelo Fernandez | TraslaIA
info@traslaia.com | https://agentcontrolprotocol.xyz
Appendix A Glossary
Term Definition
ACR-1.0 ACP Compliance Runner 1.0. Sequence-based test runner for ACP-RISK-2.0. Distinct from ACP-CR-1.0 (Change Request Process governance document).
AgentID Cryptographic identifier: base58(SHA-256(Ed25519_public_key)). Immutable and unforgeable.
Capability Token (CT) Signed JSON artifact authorizing an agent to perform specific actions on a defined resource during a limited period.
Execution Token (ET) Single-use artifact issued after an APPROVED decision. Authorizes exactly that action at that moment.
ITA Institutional Trust Anchor. Authoritative registry linking institution_id to institutional Ed25519 public key.
RIK Root Institutional Key. Institution’s Ed25519 key pair held in HSM.
Risk Score (RS) Integer in $[0,100]$ produced by the deterministic risk function. Determines the authorization decision.
Autonomy Level Integer 0–4 assigned to an agent determining applicable risk evaluation thresholds.
PoP Proof-of-Possession. Cryptographic proof that the bearer of a CT possesses the corresponding private key.
Audit Ledger Chain of signed events where $h_{n}=\text{SHA-256}(e_{n}\|h_{n-1})$. Append-only and immutable.
MRA Mutual Recognition Agreement. Bilateral document signed by two ITA authorities for cross-authority interoperability.
ESCALATED ACP decision when RS is in the intermediate range. Action not executed until explicit resolution.
Fail Closed Design principle: on any internal failure, the action is denied. Never approved by default.
Table 15: ACP Glossary.
Appendix B Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
The tla/ACP.tla module is a complete TLC-runnable TLA+ model of the ACP-RISK-2.0 admission control engine. It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$):
---- MODULE ACP ----
EXTENDS Sequences, Integers, TLC
CONSTANTS Agents, Capabilities, Resources
CapabilityBase(cap) ==
    CASE cap = "admin"     -> 60
      [] cap = "financial" -> 35
      [] cap = "write"     -> 10
      [] cap = "read"      -> 0
      [] OTHER             -> 20
ResourceScore(res) ==
    CASE res = "restricted" -> 45
      [] res = "sensitive"  -> 15
      [] res = "public"     -> 0
      [] OTHER              -> 0
ComputeRisk(cap, res) ==
    LET raw == CapabilityBase(cap) + ResourceScore(res)
    IN IF raw > 100 THEN 100 ELSE raw
Decide(rs) ==
    IF      rs >= 70 THEN "DENIED"
    ELSE IF rs >= 40 THEN "ESCALATED"
    ELSE                  "APPROVED"
VARIABLE ledger
INIT == ledger = << >>
EvaluateRequest(a, cap, res) ==
    /\ Len(ledger) < 5
    /\ ledger’ = Append(ledger,
            [ agent      |-> a,
              capability |-> cap,
              resource   |-> res,
              risk_score |-> ComputeRisk(cap, res),
              decision   |-> Decide(ComputeRisk(cap, res)) ])
[…]
            EvaluateRequest(a, cap, res)
Spec == INIT /\ [][NEXT]_ledger
Safety ==
[…]
        ledger[i].decision = "APPROVED" => ledger[i].risk_score <= 39
LedgerAppendOnly ==
[…]
        /\ ledger[i].decision $\in$ {"APPROVED", "ESCALATED", "DENIED"}
        /\ ledger[i].risk_score >= 0
RiskDeterminism ==
[…]
            ( ledger[i].capability = ledger[j].capability
           /\ ledger[i].resource   = ledger[j].resource )
            ledger[i].risk_score = ledger[j].risk_score
LedgerAppendOnlyTemporal ==
    [][ /\ Len(ledger’) >= Len(ledger)
[…]
TLC result. Running java -jar tla2tools.jar -config ACP.cfg ACP.tla with $|\mathit{Agents}|=2$, $|\mathit{Capabilities}|=4$, $|\mathit{Resources}|=3$, ledger bound =5=5 produces: Model checking completed. No error has been found. All four declared invariants/properties (TypeInvariant, Safety, LedgerAppendOnly, RiskDeterminism) and the temporal property LedgerAppendOnlyTemporal hold over the full reachable state space.
Interpretation. Safety encodes fail-closed admission: every APPROVED decision had $\mathit{RS}\leq 39$. LedgerAppendOnly encodes tamper-evidence in a state-based view: every ledger entry carries a valid decision and a non-negative risk score. LedgerAppendOnlyTemporal is the temporal statement: in every step, existing entries are preserved and the ledger never shrinks. RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot.
Acknowledgements
The ACP specification emerged from analysis of operational challenges in deploying autonomous agents in regulated B2B environments. The protocol design was informed by the IETF RFC process, the Kubernetes admission control architecture, and the SPIFFE/SPIRE workload identity model.